OpenStudy (anonymous):

please help: Use the substitution method to solve the linear system. x-y=10 5x-y=-6 and(different problem) 4x+y=2 x-y=-17

OpenStudy (anonymous):

50+5y-y=-6 4y=-56 y=-14 x-(-14)=10 x+14=10 x=-4 x=-4 and y=-14 for the first one

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

x-(2-4x)=-17 x-2+4x=-17 5x-2=-17 5x=-15 x=-3 y=14 x=3 and y=14 for the second one

OpenStudy (anonymous):

no problem :)